OpenStudy (anonymous):

Only #3 i have no clue what to do

OpenStudy (fortytherapper):

Aight

OpenStudy (usukidoll):

What is the question?

OpenStudy (anonymous):

All of question 3

OpenStudy (fortytherapper):

Which is?

OpenStudy (usukidoll):

The question is not even posted on here. Can you please type it out or provide a screenshot? "Only #3..." tells me and @FortyTheRapper nothing.

OpenStudy (anonymous):

Its in the attachment 🤔

OpenStudy (usukidoll):

that's better... so we need to use the average rate of change formula which is \[\frac{f(x)-f(a)}{x-a}\]

OpenStudy (usukidoll):

so the question is asking between 1995 and 2001 \[\frac{f(2001)-f(1995)}{2001-1995}\]

OpenStudy (anonymous):

I don't know what goes where thoo

OpenStudy (anonymous):

Dont they cancel out ?

OpenStudy (usukidoll):

not exactly. based on the chart f(2001) = 431.9 and f(1995) = 332.6 \[\frac{431.9-332.6}{2001-1995}\] now subtract 431.9-332.6 on the numerator subtract 2001-1995 on the denominator and then divide to get the result